# Physics-informed approach for exploratory Hamilton–Jacobi–Bellman equations via policy iterations

### v1

@@ 5 Discussion
In standard policy iteration (e.g., Howard’s method (Howard 1960; Kerimkulov, Šiška, and Szpruch 2020)), the policy improvement step requires a pointwise maximization, which can result in discontinuous or unstable policies in nonlinear or high-dimensional systems. By contrast, the softmax-based update in exploratory control is differentiable and admits explicit $L^{2}$-Lipschitz continuity with respect to the value function gradient. This structure enables more stable policy improvement and facilitates gradient-based learning.
@@
Our framework applies to general nonlinear stochastic systems with compact control constraints. By using physics-informed neural networks (PINNs) and mesh-free residual minimization, we avoid spatial discretization and enable tractable approximation even in high dimensions, as demonstrated on 10D LQR problems and nonlinear environments like pendulum and cartpole.
We also address some limitations and future directions.
@@
Approximation bottlenecks.
Although the global error does not accumulate theoretically, in practice, we observe that imperfect policy approximation can cause convergence plateaus. This suggests a need for stronger policy architectures, projection methods, or trust-region style updates to improve accuracy.
@@
Overall, the proposed PINN-SPI method bridges classical policy iteration theory and modern neural approximation by leveraging the analytic structure of entropy-regularized control. The result is a mesh-free, provably convergent algorithm that performs well across a broad range of benchmarks. Future work will aim to relax model assumptions, improve policy training stability, and extend to partially observed or data-driven settings.
Acknowledgement
Yeongjong Kim, Minseok Kim and Yeoneung Kim are supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government(MSIT) (RS-2023-00219980, RS-2023-00211503). The authors would like to thank Professor Hung Vinh Tran (University of Wisconsin–Madison) and Diogo Gomes (KAUST) for his insightful suggestions and valuable guidance in developing and refining the ideas of this work.
@@ Appendix A Proof of Proposition 1
where $f(x,u,p):=b(x,u)\cdot p+r(x,u).$ Since $b(x,u),r(x,u)$ are bounded and the value functions $\|v_{n}\|_{C^{2}(\mathbb{R}^{d})}$ is bounded (Tran, Wang, and Zhang 2025; Ma, Wang, and Zhang 2024), we obtain a uniform bound on the numerator:
@@ Appendix B Proof of Lemma 2
$\partial_{p_{j}}\Phi=\frac{1}{\lambda D(p)}Z(p,u)\Bigl{(}b_{j}(x,u)-\bar{b}_{j}(p)\Bigr{)}$
@@
Taking $L^{2}(U)$–norms, and recalling
$L_{\Phi}:=\frac{2\sup_{|p|\leq\tilde{L}}\Phi(p)\sqrt{d}|U|^{1/2}}{\lambda},$
@@ Appendix C Proof of Lemma 3
With $\|f\|_{H^{1}(\mathcal{X})}:=\|f\|_{L^{2}(\mathcal{X})}+\|\nabla_{x}f\|_{L^{2}(\mathcal{X})}$, we note first that $\|v^{\pi}\|_{H^{1}}+\|\tilde{v}^{\tilde{\pi}}\|_{H^{1}}<C_{1}$ by the elliptic regularity theory (Evans 2022). Set $e:=v^{\tilde{\pi}}-v^{\pi}$:
@@
where $C_{1}$ is from Assumption 1. For the entropy term, we recall Proposition 1), and set $L_{H}:=\sup_{s\in[m,M]}|1+\log s|$ where $m,M$ are from the proposition to obtain
$\|\mathcal{H}^{\pi}-\mathcal{H}^{\tilde{\pi}}\|_{2}\leq L_{H}\|\tilde{\pi}-\pi\|_{L^{2}(\mathcal{X}\times U)}$
@@
$\begin{split}&(\rho-\tfrac{1}{2}B)\|e\|_{2}^{2}+\frac{1}{2C_{0}}\|\nabla_{x}e\|_{2}^{2}\\ &\qquad\leq\underbrace{(C_{1}C_{2}+C_{1}+\lambda C_{1})L_{H}}_{=:C}\|\tilde{\pi}-\pi\|_{L^{2}(\mathcal{X}\times U)}\|e\|_{2},\end{split}$

### v2

@@ 5 Discussion
In standard policy iteration (e.g., Howard’s method (howard1960dynamic; KerimkulovSiskaSzpruch2020)), the policy improvement step requires a pointwise maximization, which can result in discontinuous or unstable policies in nonlinear or high-dimensional systems. By contrast, the softmax-based update in exploratory control is differentiable and admits explicit $L^{2}$-Lipschitz continuity with respect to the value function gradient. This structure enables more stable policy improvement and facilitates gradient-based learning.
@@
Our framework applies to general nonlinear stochastic systems with compact control constraints. By using physics-informed neural networks (PINNs) and mesh-free residual minimization, we avoid spatial discretization and enable tractable approximation even in high dimensions.
We now address some limitations and future directions.
@@
The proposed PINN-SPI method bridges classical policy iteration theory and modern neural approximation by leveraging the analytic structure of entropy-regularized control. The result is a mesh-free, provably convergent algorithm that performs well across a broad range of benchmarks. Future work will aim to relax model assumptions, improve policy training stability, and extend to partially observed or data-driven settings.
Acknowledgements
This work was supported by Seoul National University of Science and Technology. Yeoneung Kim is supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2023-00219980, RS-2023-00211503). Namkyeong Cho was supported by the Gachon University research fund of 2025 (GCU-202502800001). The authors would like to thank Professor Hung Vinh Tran (University of Wisconsin–Madison) for his insightful suggestions and valuable guidance in developing and refining the ideas of this work.
@@ Appendix A Proof of Proposition 1
where $f(x,u,p):=b(x,u)\cdot p+r(x,u).$ Since $b(x,u),r(x,u)$ are bounded and the value functions $\|v_{n}\|_{C^{2}(\mathbb{R}^{d})}$ is bounded (tran2025policy; ma2024convergence), we obtain a uniform bound on the numerator:
@@ Appendix B Proof of Lemma 2
$\partial_{p_{j}}\Phi=\frac{1}{\lambda D(p)}Z(p,u)\Bigl(b_{j}(x,u)-\bar{b}_{j}(p)\Bigr),$
@@
Taking $L^{2}(U)$norms, and recalling
$L_{\phi}:=\frac{2\|b\|_{L^{\infty}}\sup_{|p|\leq\tilde{L}}\Phi(p)\sqrt{d}|U|^{1/2}}{\lambda},$
@@ Appendix C Proof of Lemma 3
With $\|f\|_{H^{1}(\mathcal{X})}:=\|f\|_{L^{2}(\mathcal{X})}+\|\nabla_{x}f\|_{L^{2}(\mathcal{X})}$, we note first that $\|v^{\pi}\|_{H^{1}(\mathcal{X}))}+\|\tilde{v}^{\tilde{\pi}}\|_{H^{1}(\mathcal{X})}<\infty$ by the elliptic regularity theory (evans2022partial). Setting $e:=v^{\tilde{\pi}}-v^{\pi}$, we have that
@@
where $C_{1}$ is from Assumption 1. For the entropy term, we recall Proposition 1, and set $L_{H}:=\sup_{s\in[m,M]}|1+\log s|$ where $m,M$ are from the proposition to obtain
$\|\mathcal{H}^{\pi}-\mathcal{H}^{\tilde{\pi}}\|_{2}\leq L_{H}\|\tilde{\pi}-\pi\|_{L^{2}(\mathcal{X}\times U)}.$
@@
$\begin{split}&(\rho-\tfrac{1}{2}B)\|e\|_{2}^{2}+\frac{1}{2C_{0}}\|\nabla_{x}e\|_{2}^{2}\\ &\qquad\leq\underbrace{(C_{1}C_{2}+C_{1}+\lambda C_{1})L_{H}}_{=:C}\|\tilde{\pi}-\pi\|_{L^{2}(\mathcal{X}\times U)}\|e\|_{2}.\end{split}$
@@
Appendix D Experimental details
D.1 Stochastic linear-quadratic regulator.
For the stochastic LQR problem where the dynamics is given by
$\mathrm{d}X_{t}=(AX_{t}+Bu_{t})\mathrm{d}t+\sigma\mathrm{d}W_{t},$
we use
$A=\begin{bmatrix}0.0446&0.0300&0.4497&0.2813&0.0311\\ 0.0432&0.2885&0.0321&0.0252&0.0419\\ 0.1975&0.0430&0.0301&0.0422&0.2282\\ 0.0436&0.4716&0.3812&0.3010&0.0372\\ 0.0225&0.0629&0.0335&0.0424&0.1836\\ \end{bmatrix}$,
$B=\begin{bmatrix}0.0681&0.0544&0.0467&0.0152&0.0787\\ 0.0970&0.0081&0.0145&0.0034&0.0984\\ 0.0358&0.0833&0.0324&0.0839&0.0012\\ 0.0116&0.0280&0.0056&0.0092&0.0432\\ 0.0047&0.0848&0.0718&0.0977&0.0556\\ \end{bmatrix}$
for $d=m=5$, and
$A=\begin{bmatrix}0.0866&0.0191&0.1394&0.1392&0.1480&0.0241&0.0988&0.1435&0.1883&0.1923\\ 0.1271&0.0900&0.0924&0.0766&0.0696&0.0466&0.1716&0.0371&0.1097&0.0946\\ 0.0595&0.1656&0.1953&0.1353&0.1872&0.0587&0.0830&0.0035&0.0215&0.0740\\ 0.1971&0.0808&0.1301&0.0157&0.1908&0.1505&0.0662&0.1334&0.1394&0.1951\\ 0.0570&0.0419&0.0470&0.0916&0.1094&0.0640&0.0159&0.1687&0.1224&0.0294\\ 0.1137&0.1033&0.0379&0.0881&0.1224&0.0139&0.0060&0.1857&0.0732&0.0989\\ 0.1271&0.0414&0.1232&0.1896&0.1457&0.0997&0.1830&0.1309&0.0673&0.0855\\ 0.0827&0.1076&0.1498&0.1164&0.0192&0.1888&0.1357&0.1352&0.1086&0.1959\\ 0.1489&0.0223&0.0018&0.0002&0.1631&0.1272&0.0282&0.0075&0.0351&0.0478\\ 0.0791&0.1215&0.0219&0.1653&0.0635&0.0230&0.1943&0.0373&0.0253&0.1129\end{bmatrix}$,
$B=\begin{bmatrix}0.0728&0.0209&0.0336&0.0269&0.0197&0.0844&0.0358&0.0483&0.0033&0.0559\\ 0.0459&0.0812&0.0411&0.0338&0.0581&0.0689&0.0084&0.0015&0.0198&0.0053\\ 0.0136&0.0803&0.0666&0.0570&0.0277&0.0721&0.0613&0.0106&0.0115&0.0699\\ 0.0848&0.0815&0.0915&0.0766&0.0683&0.0997&0.0389&0.0485&0.0630&0.0102\\ 0.0858&0.0205&0.0769&0.0968&0.0722&0.0004&0.0201&0.0990&0.0836&0.0750\\ 0.0168&0.0033&0.0286&0.0740&0.0314&0.0238&0.0183&0.0277&0.0889&0.0123\\ 0.0795&0.0445&0.0037&0.0776&0.0038&0.0103&0.0183&0.0542&0.0722&0.0544\\ 0.0399&0.0139&0.0071&0.0227&0.0556&0.0885&0.0062&0.0271&0.0382&0.0641\\ 0.0142&0.0063&0.0456&0.0536&0.0993&0.0206&0.0264&0.0463&0.0695&0.0907\\ 0.0301&0.0514&0.0583&0.0007&0.0900&0.0426&0.0385&0.0077&0.0110&0.0930\end{bmatrix}$
for $d=m=10$. The reward function was set to
$L(x,u)=-x^{\top}Qx-u^{\top}Ru,$
with compact action set $U=\{u\in\mathbb{R}^{m}:\|u\|_{\infty}\leq\bar{u}\}$. We set $Q=5I_{d}$, $R=I_{m}$, $\bar{u}=10$, and use isotropic noise $\sigma=0.1I_{d}$. The entropy weight is fixed to $\lambda=0.1$. The total return is computed as the mean over 30 trajectories with time step $\Delta t=0.01$, horizon $T=2$, and discount rate $\rho=1$.
PINN–SPI.
For all dimensions $d\in\{5,10,20\}$, we use the same network architecture and hyperparameters. The value network $v(x;\theta)$ is an MLP with hidden width $100$ and depth $2$, SiLU activations, and residual connections between hidden layers. It is optimized by Adam with learning rate $3\times 10^{-4}$. The policy network $\pi(x,u;\omega)$ is an MLP with hidden width $100$ and depth $4$, also using SiLU and residual connections, trained by Adam with learning rate $3\times 10^{-5}$. At each policy-iteration step we draw $100$ state collocation points and, for every state, approximate the entropy terms using $100$ action samples on the compact control set. We perform $1000$ outer policy-iteration steps, and in each step run $10$ epochs of value-network training and $10$ epochs of policy-network training.
SAC.
For SAC, both the actor and critic are MLPs with hidden width $256$, depth $3$, ReLU activations, and output layers adapted to the action space. All networks are trained with Adam with learning rate $1\times 10^{-5}$. We use a replay buffer with mini-batch size $256$, Polyak averaging coefficient $\tau=0.005$, and entropy regularization coefficient $0.2$. Training is performed over $2000$ gradient-update iterations, with $500$ warm-up steps before starting policy updates.
PPO.
For PPO, the actor and critic share the same MLP architecture: hidden width $256$, depth $3$, ReLU activations, trained with Adam. We use clipping parameter $0.2$, generalized advantage estimation $\lambda_{\mathrm{GAE}}=0.95$, value loss coefficient $0.5$, entropy coefficient $0.01$, and gradient clipping with $\texttt{max\_grad\_norm}=1.0$. For the 55D LQR problem, we set the actor and critic learning rates to $3\times 10^{-5}$, use $4096$ rollout steps per update, $10$ PPO epochs, and mini-batch size $256$. For the 1010D problem, the actor and critic learning rates are $2\times 10^{-5}$ and $6\times 10^{-5}$, respectively, with $1024$ rollout steps, $6$ epochs, and mini-batch size $512$.
D.2 Stochastic Cartpole
We consider the stochastic cartpole environment with discount factor $\rho=0.5$ and entropy temperature $\lambda=0.1$. The total reward is computed as the average over $10$ trajectories.
PINN–SPI.
Since the action space is discrete, we do not parameterize a separate policy network. Instead, we apply the softmax update in (4) and act greedily by selecting the action with the highest reward. The value function $v(x;\theta)$ is represented by an MLP with hidden width $100$, depth $2$, SiLU activations, and residual connections, trained by Adam with learning rate $3\times 10^{-4}$. PINN–SPI is run for $30$ outer iterations; in each iteration we perform $3000$ gradient steps using $10000$ state collocation points. For each state we evaluate all available actions, which is equivalent to using one action sample per state.
SAC.
The SAC actor and critic are MLPs with hidden width $256$, depth $3$, ReLU activations, trained with Adam with learning rate $3\times 10^{-4}$. We use mini-batch size $256$, Polyak update coefficient $\tau=0.005$, entropy coefficient $0.2$, and $300$ warm-up steps. Training is run for $200$ iterations.
PPO.
The PPO actor and critic are again $256$-wide, 33-layer MLPs with ReLU activations and trained with Adam optimizer with learning rate $10^{-4}$. We use $1024$ rollout steps per update, mini-batch size $256$, $10$ PPO epochs, clipping parameter $0.2$, GAE parameter $\lambda_{\mathrm{GAE}}=0.95$, value loss coefficient $0.5$, entropy coefficient $0.01$, and maximum gradient norm $1.0$.
Stochastic Pendulum
For the stochastic pendulum, we use discount factor $\rho=0.5$ and report the average total reward over $10$ trajectories.
PINN–SPI.
The value network $v(x;\theta)$ is an MLP with hidden width $100$ and depth $2$, using SiLU activations and residual connections, trained by Adam with learning rate $3\times 10^{-4}$. The policy network $\pi(x,u;\omega)$ has hidden width $100$ and depth $4$, also with SiLU and residual connections, trained with Adam with learning rate $3\times 10^{-5}$. PINN–SPI is executed for $12$ outer iterations; in each iteration we perform $1000$ gradient steps for both value and policy networks. We use $100$ state collocation points for the value network and $1000$ state–action collocation points for the policy network, corresponding to $1000$ action samples per state on the continuous action domain. The entropy weight is fixed to $\lambda=0.1$.
SAC.
The SAC actor and critic are $256$-wide, 3-layer MLPs with ReLU activations, trained with Adam with learning rate $3\times 10^{-4}$. We use a mini-batch size of $256$, entropy coefficient $0.2$, Polyak coefficient $\tau=0.005$, and run 20 00020\,000 gradient iterations.
PPO.
For PPO, the actor and critic are again $256$-wide, 3-layer MLPs with ReLU activations, trained with Adam with learning rate $10^{-4}$ for both networks. We use $1024$ rollout steps, mini-batch size $256$, $10$ PPO epochs per update, clipping parameter $0.2$, GAE parameter $\lambda_{\mathrm{GAE}}=0.95$, value loss coefficient $0.5$, entropy coefficient $0.01$, and maximum gradient norm $1.0$.
Appendix E Additional experimental result
E.1 20D stochastic LQR
Consistent with the 5D and 10D LQR settings, the 20D stochastic LQR experiment also exhibits a clear upward trend in the cumulative reward as training progresses, in agreement with our theoretical convergence results. This demonstrates the scalability of our proposed approach to higher-dimensional stochastic control problems.
For the 20D system, the matrices $A$ and $B$ are generated by sampling each entry independently from the uniform distribution on $[0,0.5]$, following the same protocol used in the 5D and 10D experiments. We employ identical experimental configurations for PINN–SPI and SAC as in the lower-dimensional cases. For PPO, we adopt a slightly modified setup: both the actor and critic learning rates are set to $3\times 10^{-5}$, and we use $512$ rollout steps per update, $4$ training epochs, and a mini-batch size of $256$.
Figure 4 illustrates that our method yields competitive performance and consistently attains higher cumulative rewards throughout training.
Figure 4: Comparison between our method and others, SAC and PPO.
E.2 Verification of monotonicity
A key property of soft policy iteration (tran2025policy) is the monotonic improvement of the value function under exact updates. Our PINN–SPI method numerically preserves this monotonicity despite the presence of approximation errors from neural representation and finite collocation as shown in Figure 5
(a) Cartpole: Stable and monotonic convergence.
(b) Pendulum: PINN-SPI reaches high total reward.
Figure 5: Evaluation reward over training time for PINN-SPI on cartpole and cendulum tasks. The average total reward tends to increase as policy iteration proceeds.